News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 October 2001 12.00pm EDT

During the last week, Chandra completed the observing schedule as planned.

A real-time procedure was completed on Oct 16 that verified the CPE warm start count has remained unchanged at 1 since Mar 2. A second procedure was completed on Oct 18 to dump and verify values of the gyroscope scale factors and related variables. The dump and verification was done in preparation for an update to the scale factors.

The schedule of observations for the coming week is shown below.

------------------------------------------
PSS 0902+1640          ACIS-S        Oct 21
PSR B0833-45           HRC-I
Radiation Belts                      Oct 22
1RXS J1708-4009        HRC-I         Oct 23
NGC 7130               ACIS-S
PSR1813-36             ACIS-S
SAX J1806.5-2215       ACIS-S        Oct 24
IRAS 17208-0014        ACIS-S
CHVC125                ACIS-S
SDSS094108+594725      ACIS-S        Oct 25
Radiation Belts
MG J0414+0534          ACIS-S
HS0218+3707            ACIS-S        Oct 26
PSS0134+3307           ACIS-S
2MASSJ222202+1952      ACIS-I
M17                    ACIS-I
SAX J1747.0-2853       ACIS-S/HETG
SN1998BW               ACIS-S        Oct 27

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports